9. Working with Numbers and Graphs Q9 The following table summarizes the combinations of goods X and Y that two individuals, Tina and David, can each produce individually. Tina David 250 02000 125 62.5 100 10 0 125 0 20 If each person specializes in the production of the good for which they have the comparative advantage, then production of good Y, an individual.) ▼ would specialize in the d would specialize in the production of good X. (Hint: Calculate the opportunity cost of producing X and Y for each Suppose that David and Tina each specializes in the production of the good in which each has a comparative advantage. Fill in the top row of the following table to indicate Davids and Tinas output of each good. Tina David Before Trade, with Specialization After Trade After specialization, suppose the two agree on a price of 3 units of good X for each unit of good Y. When the two individuals make the trade, they exchange 150 units of good X for 50 units of good Y. Recall that the individual who has specialized in the production of X would trade 150 units of good X and receive 50 units of good Y, while the individual who specialized in the production of good Y would trade 50 units of good Y and receive 150 units of good X. In the previous table, fill in the second row to indicate the amounts of good X and good Y each individual has after the trade. The following graphs plot Tinas production possibilities frontier (PPF) in purple and Davids PPF in blue, respectively. Using the second row of the previous table, plot Tinas new combination of Xand Ywith the grey point (star symbol)

Using the second row of the previous table, plot Tinas new combination of X and Y with the grey point (star symbol). 150 xX 135 Tina after Trade 120 105 90 30 15 Tinas PPP 1 82 123 164 205 246 287 328 369410 Using the second row of the previous table, plot Davids new combination of X and Y with the tan point (dash symbol) David 72 David after Trade 64 56 40 Davids PPF 16 0 244 72 98 120 44 168 12 218 240 True or False: Both individuals can now consume a combination of X and Y that was previously unattainable.

a) True

b) False

Note: This whole problem is just one question but consists of many parts. I would highly, extremely appreciate it if someone could solve this whole thing for me, this is tricky.
Answer #1

Tina:

250 X = 125 Y

X = 125/250 = 0.5 Y

Y = 250/125 = 2 X

Opportunity cost of 1 X is 0.5 Y and 1 Y is 2 X.

David:

200 X = 20 Y

1 X = 20/200 = 0.1 Y

1 Y = 200/20 = 10 X

Opportunity cost of 1 X is 0.1 Y and 1 Y is 10 X.

Since, opportunity cost of producing X is lower when it is produced by David. So, David has comparative advantage in the production of X and Tina has comparative advantage in the production of Y.

20000 Before Trade, with Specialization After Trade 12% 150 | 125-50+75 1200-150-50 50

David Tina 150 72 135 Tina after Trade 64 120 58 105 48 40 60 2 Davids PPF 45 16 30 15 Tinas PPF 0 24 4 72 98 20 144 168 192 216 240 0 41 82 123 164 205 246 287 328 369 410A) True

Before trade, person can consume only bundles which lies below or on PPF.